OpenStudy (anonymous):

pleas help ASAP simplify the expression√10-4 √10

OpenStudy (tkhunny):

Can you simplify this...? R - 4R = ??

OpenStudy (anonymous):

i think i can why

OpenStudy (tkhunny):

Why is it different from the problem you posted?

OpenStudy (anonymous):

cause i am tryna help my sis and i am not soo good in math

OpenStudy (tkhunny):

You didn't answer my question. \(R - 4R = -3R\) \(\sqrt{10} - 4\sqrt{10} = \)??

OpenStudy (anonymous):

when I did \[\sqrt{10-4} \sqrt{10}\] I had gotten 2.44 and 3.16

OpenStudy (tkhunny):

Is that the real problem? That was not clear in the original post. \(\sqrt{10-4} = \sqrt{6} = \sqrt{2\cdot 3} = \sqrt{2}\cdot\sqrt{3}\) \(\sqrt{10} = \sqrt{2\cdot 5} = \sqrt{2}\cdot\sqrt{5}\) There are various ways to play with it. None, really, to "simplify" it.

OpenStudy (anonymous):

okay here are the answers i have a 5√10 b -3√10 c -3√20 d 5√20

OpenStudy (tkhunny):

Okay, then I repeat: You didn't answer my question. \(R−4R=−3R\) \(\sqrt{10}−4\sqrt{10} = \)?? \(Dog−4Dog=−3Dog\) \(\sqrt{10}−4\sqrt{10} = \)??

OpenStudy (anonymous):

is the answer 5√10 then